OpenStudy (anonymous):

I have the following integral:

OpenStudy (anonymous):

\[I_{n} = \int\limits_{0}^{1}\frac{ x^{n} }{ 1+x } dx\]

OpenStudy (anonymous):

and i also have to prove that:

OpenStudy (shamim):

ok i wanna solve it

OpenStudy (anonymous):

\[n I_{n} = \frac{ 1 }{ 2 } - \int\limits_{0}^{1} \frac{ x^{n} }{ (1+x)^{2} }\]

OpenStudy (anonymous):

and dx of course. :)

OpenStudy (anonymous):

i think it looked better when you used integration by parts, but we couldn t get \[x^{n}\]at the numerator.

OpenStudy (anonymous):

i was thinking about trying to write \[\frac{ 1 }{ 2 }\] as \[\int\limits_{0}^{1} \] from something and then try to bring it to the final form.